Bryant Company has a factory machine with a book value of $93,700 and a remaining useful life of 7 years. It can be sold for $34,700. A new machine is available at a cost of $378,500. This machine will have a 7-year useful life with no salvage value. The new machine will lower annual variable manufacturing costs from $605,900 to $457,900. Prepare an analysis showing whether the old machine should be retained or replaced. (In the first two columns, enter costs and expenses as positive amounts, and any amounts received as negative amounts. In the third column, enter net income increases as positive amounts and decreases as negative amounts. Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)
Answer #1
Retain Equipment Replace Equipment Net income increase(decrease)
Variable manufacturing costs 4241300 3205300 1036000
New machine cost 0 378500 -378500
Sell old machine 0 -34700 34700
Total 4241300 3549100 692200
The old factory machine should be replaced
Workings:
Retain Equipment Replace Equipment
Variable manufacturing costs =605900*7 =457900*7